# BianCang: A Traditional Chinese Medicine Large Language Model

### v1

@@ Abstract
The rise of large language models (LLMs) has driven significant progress in medical applications, including traditional Chinese medicine (TCM). However, current medical LLMs struggle with TCM diagnosis and syndrome differentiation due to substantial differences between TCM and modern medical theory, and the scarcity of specialized, high-quality corpora. This paper addresses these challenges by proposing BianCang (扁仓), a TCM-specific LLM, using a two-stage training process that first injects domain-specific knowledge and then aligns it through targeted stimulation. To enhance diagnostic and differentiation capabilities, we constructed pre-training corpora, instruction-aligned datasets based on real hospital records, and the ChP-TCM dataset derived from the Pharmacopoeia of the People’s Republic of China. We compiled extensive TCM and medical corpora for continuous pre-training and supervised fine-tuning, building a comprehensive dataset to refine the model’s understanding of TCM. Evaluations across 11 test sets involving 29 models and 4 tasks demonstrate the effectiveness of BianCang, offering valuable insights for future research. Code, datasets, and models are available at https://github.com/QLU-NLP/BianCang.
@@ 1 Introduction
Recently, LLMs have achieved remarkable success in natural language processing [Zhao et al., 2023]. General-purpose models such as GPT-4 [Achiam et al., 2023], Qwen2 [Yang et al., 2024a], and Baichuan2 [Yang et al., 2023] have demonstrated impressive capabilities across diverse tasks. Building on these models, researchers have effectively applied LLMs to specialized domains such as healthcare [He et al., 2023], law [Yue et al., 2023], and finance [Chen et al., 2023b], often at relatively low cost. Among these fields, healthcare, stands out due to its critical role in safeguarding human health and extending life, making it a focal point of extensive research and practical applications.
Figure 1: Different diagnostic processes of TCM and modern medicine for the same medical record.
The complexity and specialization of knowledge in the biomedical field place higher demands on the domain expertise of LLMs [He et al., 2023]. To address this challenge, secondary training on general-purpose LLMs is often required. Researchers typically focus on optimizing the training methods and improving the quality of training data to enhance the effectiveness of this process. From the perspective of optimizing training methods, Zhang et al. introduced reinforcement learning from mixed feedback (RLMF) to integrate data from various sources, resulting in the development of HuatuoGPT1 [Zhang et al., 2023a]. Chen et al. proposed a training method combining pre-training and supervised fine-tuning into a single stage, simplifying the process and leading to HuatuoGPT2 [Chen et al., 2023a]. To improve the quality of training data, Bao et al. employed diverse strategies to construct the DISC-Med-SFT dataset, which enhanced the responsiveness of medical LLMs in both single-round and multi-round consultation scenarios, resulting in DISC-MedLLM [Bao et al., 2023]. Luo et al. developed a multi-task bilingual Chinese-English dataset to train the Taiyi model [Luo et al., 2024a].
@@
These medical LLMs perform well in modern medicine, but their performance in TCM is relatively unsatisfactory. The primary reason for this is the significant difference between TCM theory and modern medicine. TCM has a unique and complete theoretical foundation. Unlike modern medicine, which treats diseases based on their types, TCM determines the patient’s syndrome type by analyzing the evidence collected through the “four diagnostic methods” (望闻问切: inspection, listening and smelling, inquiry, and palpation). Treatment is then administered based on the syndrome type. As a result, patients with the same disease and different syndrome types may receive different treatments, while patients with different diseases and the same syndrome type may receive the same treatment. This approach, known as “treating the same disease with different therapies” (同病异治) and “treating different diseases with the same therapy” (异病同治), is central to TCM [Mucheng et al., 2022].
@@
As shown in Fig. 1, for the same patient, modern medicine uses medical instruments to measure the T-wave (related to changes in the electrocardiogram) and diagnoses the patient based on the specific value changes and trends of the T-wave. In contrast, TCM maps the patient’s described symptoms and the evidence gathered from the four diagnostic methods into a unique Yin-Yang (阴-阳) coordinate system, analyzes the underlying causes, and synthesizes these findings to determine the patient’s syndrome type. Compared to modern medicine, which relies on measurable data from instruments, TCM is more abstract, experience-based, and harder to explain [Mucheng et al., 2022]. Facing the complexity of TCM tasks and the limitations of LLMs, Zhang et al. trained Qibo using a specialized TCM corpus and explored applying Qibo to TCM consultations [Zhang et al., 2024]. Yang et al. trained Zhongjing using a self-built dataset of authentic doctor-patient dialogues, aiming to improve the LLM’s ability to conduct proactive inquiries and understand multi-round consultations in TCM [Yang et al., 2024b]. Hua et al. focused on enhancing the model’s grasp of TCM and pharmacology, training Lingdan [Hua et al., 2024]. While these TCM-specific LLMs have made some progress, their ability to differentiate and diagnose syndromes in TCM remains limited, particularly in real diagnostic scenarios where practical syndrome differentiation and diagnostic analysis skills are required.
@@
To address the aforementioned challenges and further enhance the comprehensive capabilities of TCM-focused LLMs, we propose BianCang (扁仓), a large TCM language model based on Qwen-2/2.5, designed to improve syndrome differentiation and diagnostic capabilities in TCM. BianCang adopts a two-stage training process: in the first stage, extensive traditional Chinese medicine and medical knowledge is injected through continuous pre-training of the foundational model; in the second stage, supervised fine-tuning is applied to activate and align the model’s internal knowledge. We evaluated BianCang on 11 test sets across 4 tasks, and the results show that it consistently outperforms existing open-source TCM-specific and Chinese medical language models in all dimensions of capability.
@@
  • We developed BianCang, a novel TCM LLM, through a two-stage training process tailored to TCM characteristics. The model has acquired extensive TCM knowledge and real clinical case expertise, demonstrating genuine syndrome differentiation and diagnostic capabilities.
  • We constructed and open-sourced the ChP-TCM dataset, based on the Pharmacopoeia of the People’s Republic of China, designed for pre-training and supervised fine-tuning of TCM models.
  • We performed a comprehensive evaluation of existing medical and TCM LLMs from multiple dimensions, considering both objective and subjective aspects. Extensive experiments involving 11 test sets, 29 models, and 4 tasks validate the effectiveness and robustness of the BianCang model, providing valuable experimental comparisons for future research.
@@ 2.1 LLMs in Chinese Medical Domain
Among the existing Chinese medical LLMs, some works are based on open-source foundation models, which are then fine-tuned on self-constructed medical datasets. Examples include DoctorGLM [Xiong et al., 2023] and BenTsao [Wang et al., 2023]. To further enhance the capabilities of medical LLMs, researchers have explored optimizing training methods from various perspectives. For instance, Zhang et al. introduced RLMF (Reinforcement Learning from Mixed Feedback) in their training process to coordinate both ChatGPT-extracted data and real-world data from doctors, resulting in HuatuoGPT1 [Zhang et al., 2023a]. Chen et al., aiming to reduce the complexity of multi-stage training strategies, proposed a one-stage domain adaptation protocol where heterogeneous data from both traditional pre-training and supervised stages are unified into a simple instruction-output pair format, which they used to train HuatuoGPT2 [Chen et al., 2023a]. Tian et al. developed ChiMedGPT, following a process that included pre-training, supervised fine-tuning (SFT), and reinforcement learning from human feedback (RLHF) [Tian et al., 2024]. Other researchers have attempted to improve the quality of training data. For example, Bao et al. constructed a high-quality SFT dataset using three strategies: utilizing medical knowledge graphs, reconstructing real-world dialogues, and incorporating human-guided preference rephrasing, significantly enhancing the model’s response capabilities in both single-turn and multi-turn consultation scenarios [Bao et al., 2023]. Luo et al. built a rich, multi-task bilingual dataset in both Chinese and English and trained the Taiyi model [Luo et al., 2024a]. Although these medical LLMs have achieved considerable success in modern medicine, their performance in the TCM domain is significantly lacking.
@@ 2.2 LLMs in TCM Domain
Previous Chinese medical LLMs have mostly focused on modern medical systems. Unlike modern medicine, TCM emphasizes a holistic view and individualized treatment, focusing on balancing Yin-Yang, harmonizing Qi (气) and blood (血), and coordinating the internal organs. TCM diagnoses are conducted through the four diagnostic methods (inspection, listening/smelling, inquiry, and palpation) to understand the patient’s overall conditions [Mucheng et al., 2022]. The theoretical foundation of syndrome differentiation and disease diagnosis in TCM is unique and fundamentally different from modern medicine. Existing medical LLMs are unable to grasp the principles of TCM, resulting in unsatisfied performance.
Some studies have specifically aimed at improving the capabilities of LLMs in TCM. For example, Zhang et al. built a TCM-specific corpus and trained the Qibo model, which was applied to the pipeline of TCM consultations [Zhang et al., 2024]. Yang et al. constructed a real-world, multi-turn Chinese medical dialogue dataset, CMtMedQA, and trained the Zhongjing model using continuous pre-training, SFT, and RLHF, which enhanced the model’s ability in active inquiry and multi-turn understanding within the context of TCM [Yang et al., 2024b]. Hua et al. focused on traditional Chinese medicine, constructing datasets for TCM pre-training, traditional Chinese patent medicine (TCPM) Q&A, and the spleen and stomach herbal prescription recommendation, and used these datasets to train the Lingdan model [Hua et al., 2024]. Although these efforts have made significant progress in TCM LLMs, these models still lack real-world diagnostic experience in TCM and perform poorly on real-world syndrome differentiation and disease diagnosis tasks.
Figure 2: The overall flowchart of constructing BianCang. In the first stage, extensive traditional Chinese medicine and medical knowledge is injected into the foundational model through continuous pre-training. In the second stage, supervised fine-tuning is applied to activate and align the internal knowledge of the model.
To address the aforementioned challenges, we propose BianCang, a novel TCM LLM based on Qwen2/2.5. BianCang first undergoes continuous pre-training to inject large amounts of TCM and medical knowledge, as well as real medical records, into the model. It is then fine-tuned using a carefully curated set of TCM-specific instructions to activate the model’s internal knowledge and align these instructions with the model’s internal knowledge. This training strategy ensures consistency in the model’s knowledge before and after SFT, significantly improving its performance in tasks such as syndrome differentiation and disease diagnosis in TCM.
@@ 3 Methods
This section introduces the construction of BianCang, which consists of two stages: continuous pre-training and supervised fine-tuning. The overall process is shown in Fig. 2.
@@
Table 1: The statistics of the continuous pre-training data.
3.1 Continuous Pre-training
@@
SFT is a key stage in building large language models. Recent research has found that SFT is not a process of learning world knowledge specific to a domain, but rather a process of aligning instructions with the knowledge already embedded in the parameters of LLMs [Ren et al., 2024]. In order to maintain consistency in the model’s knowledge before and after SFT as much as possible, we first perform continuous pre-training on the base model to inject extensive knowledge of traditional Chinese medicine and healthcare [Zhou et al., 2023].
Books are the crystallization of human wisdom and serve as high-quality data for training large language models [Gunasekar et al., 2023]. We collected a vast amount of TCM and medical books and processed them through OCR, cleaning, and segmentation. Similarly, we also collected open-source medical encyclopedias and literature [Chen et al., 2023a]. To further enhance the model’s knowledge of TCM, we incorporated more specialized TCM corpora into the pre-training data, including the Pharmacopoeia of the People’s Republic of China (Part I: Traditional Chinese Medicine) and common TCM syndrome knowledge bases, which we scraped and cleaned using web crawlers, as well as the ChatMed-TCM dataset [Wei et al., 2023]. For TCM, relying solely on theoretical knowledge is insufficient; practical diagnostic experience is also required. Therefore, we included 12,259 real patient case records and 56,928 real TCM syndrome differentiation and diagnosis records in the pre-training stage. Additionally, to improve the model’s performance in TCM and medical-related exams, we introduced the training set from the Chinese Medical Benchmark (CMB) [Wang et al., 2024].
@@
Besides TCM and medical corpora, we also incorporated general domain corpora into the pre-training data to prevent the model from overfocusing on special domains, which could lead to catastrophic forgetting in general domains. Among them, COIG-CQIA is a high-quality, multi-domain instruction fine-tuning dataset aligned with human interaction behaviors [Bai et al., 2024]. APE-210K is a diverse collection of mathematical word problems [Zhao et al., 2020]. Evol-Instruction-66K is a filtered and curated dataset for code generation [Luo et al., 2024b]. We concatenated the instructions and responses from these three datasets as the pre-training data. Notably, all three general-domain datasets exhibit a certain level of logical reasoning, which we speculate could benefit tasks such as TCM syndrome differentiation, as it also requires logical reasoning skills.
@@
The data statistics for pre-training are shown in Table 1. Using this data, we performed continuous pre-training on Qwen2-7B and Qwen2.5-7B [Yang et al., 2024a], eventually obtaining the BianCang TCM base model.
3.2 Supervised Fine-Tuning (SFT)

### v2

@@ Abstract
The surge of large language models (LLMs) has driven significant progress in medical applications, including traditional Chinese medicine (TCM). However, current medical LLMs struggle with TCM diagnosis and syndrome differentiation due to substantial differences between TCM and modern medical theory, and the scarcity of specialized, high-quality corpora. To this end, in this paper we propose BianCang (扁仓) [^1] The term BianCang (扁仓) is formed by combining the first characters of the names of two renowned ancient Chinese TCM physicians, Bian Que (扁鹊) and Cang Gong (仓公). , a TCM-specific LLM, using a two-stage training process that first injects domain-specific knowledge and then aligns it through targeted stimulation to enhance diagnostic and differentiation capabilities. Specifically, we constructed pre-training corpora, instruction-aligned datasets based on real hospital records, and the ChP-TCM dataset derived from the Pharmacopoeia of the People’s Republic of China. We compiled extensive TCM and medical corpora for continual pre-training and supervised fine-tuning, building a comprehensive dataset to refine the model’s understanding of TCM. Evaluations across 11 test sets involving 31 models and 4 tasks demonstrate the effectiveness of BianCang, offering valuable insights for future research. Code, datasets, and models are available on GitHub. [^2] https://github.com/QLU-NLP/BianCang
{IEEEkeywords}
Large Language Model, Natural Language Generation, Traditional Chinese Medicine
Figure 1: Different diagnostic processes of TCM and modern medicine for the same medical record. Modern medicine relies on the chief complaint and auxiliary examination on the T-wave through electrocardiograms, diagnosing the patient based on specific value changes and trends. In contrast, TCM interprets the patient’s chief complaint and diagnostic information from the four diagnostic methods within a unique Yin-Yang framework, identifying underlying causes and synthesizing the findings to determine the syndrome type. While modern medicine depends on quantifiable data, TCM is more abstract and experience-based.
@@ 1 Introduction
Recently, LLMs have achieved remarkable success in natural language processing [1, 2]. General-purpose models such as DeepSeek-V3 [3], DeepSeek-R1 [4], GPT-4 [5] and Qwen2.5 [6] have demonstrated fascinating capabilities across diverse tasks. Building on these models, researchers have effectively applied LLMs to vertical domains such as healthcare [7, 8, 9], law [10], and finance [11], often at relatively low cost by prompting, in-context learning, or continual training. Among verticalizations, healthcare stands out due to its critical role in improving human health and enhancing longevity, attracting increasing interest from both academia and industry [12, 13].
The major challenge to adapting general LLMs into the biomedical domain is a lack of accurate knowledge and expertise [14], which is critical to tackle a variety of domain-specific tasks. In general, works addressed this challenge by leveraging continual/secondary training methods through reinforcement learning and streamlined fine-tuning [15, 16] and improving data quality with specialized datasets [17, 18].
@@
Despite successful applications of the resulting medical LLMs in modern medicine, the performance of TCM still lags far behind, mainly due to theoretical differences between TCM and modern medicine, leading to further knowledge and domain gaps. To be specific, TCM has a unique and complete theoretical foundation. Unlike modern medicine, which treats diseases based on their types, TCM determines the patient’s syndrome type by analyzing the evidence collected through the “four diagnostic methods” (望/闻/问/切: inspection, listening and smelling, inquiry, and palpation). Treatment is then administered based on the syndrome type. As a result, patients with the same disease and different syndrome types may receive different treatments, while patients with different diseases and the same syndrome type may receive the same treatment. This approach, known as “treating the same disease with different therapies” (同病异治) and “treating different diseases with the same therapy” (异病同治), is central to TCM [19].
@@
Take Fig. 1 as an example: For the same patient, modern medicine uses medical instruments to measure the T-wave (related to changes in the electrocardiogram) and diagnoses the patient based on the specific value changes and trends of the T-wave. In contrast, TCM maps the patient’s described symptoms and the evidence gathered from the four diagnostic methods into a unique Yin-Yang (阴-阳) coordinate system, analyzes the underlying causes, and synthesizes these findings to determine the patient’s syndrome type. As such, TCM applications are more abstract, experience-based, and hard to explain [19], differing from modern medicine relying on measurable data from instruments.
Recently, works have explored the development of TCM-specific LLMs, including Qibo [20], which was trained on a specialized TCM corpus for TCM consultations, Zhongjing [21], which focused on doctor-patient dialogues to improve proactive inquiry and consultation understanding, and Lingdan [22], which aimed to enhance the model’s grasp of TCM and pharmacology.
However, these models still face significant challenges in practical syndrome differentiation and diagnostic analysis, where the complex and abstract nature of TCM theories and the scarcity of high-quality, structured diagnostic data hinder their ability to perform accurate and consistent syndrome differentiation in real-world diagnostic scenarios.
@@
To address the challenge above, we propose BianCang (扁仓), a large TCM language model built on Qwen-2/2.5, designed to enhance syndrome differentiation and diagnostic capabilities in TCM. BianCang employs a two-stage training process: in the first stage, extensive traditional Chinese medicine and medical knowledge are integrated through continual pre-training of the foundational model, whereas in the second stage, supervised fine-tuning is conducted to activate and align the model’s internal knowledge. We evaluated BianCang on 11 benchmark datasets across 4 tasks, and the results demonstrate that it consistently outperforms existing open-source TCM-specific and Chinese medical language models across all key performance dimensions.
Figure 2: The overall flowchart of constructing BianCang. In the first stage, extensive traditional Chinese medicine and medical knowledge is injected into the foundational model through continual pre-training. In the second stage, supervised fine-tuning is applied to activate and align the internal knowledge of the model.
@@
  • We presented BianCang, a new TCM LLM with a two-stage training process tailored to the distinctive characteristics of TCM. BianCang has acquired comprehensive TCM knowledge and real clinical expertise, demonstrating strong and consistent capabilities in syndrome differentiation and disease diagnosis.
  • We constructed and open-sourced the ChP-TCM dataset, based on the Pharmacopoeia of the People’s Republic of China, specifically designed for pre-training and supervised fine-tuning of TCM models.
  • We performed a comprehensive evaluation of existing medical and TCM LLMs from multiple perspectives, considering both objective and subjective aspects. Extensive experiments involving 11 datasets, 31 models, and 4 tasks validate the effectiveness and robustness of the BianCang model, providing valuable benchmarks for future research.
@@ 2.1 LLMs in Chinese Medical Domain
Among the existing Chinese medical LLMs, some works are based on open-source foundation models, which are then fine-tuned on self-constructed medical datasets. Examples include DoctorGLM [23] and BenTsao [15]. To further enhance the capabilities of medical LLMs, researchers have explored optimizing training methods from various perspectives. For instance, Zhang et al. introduced RLMF (Reinforcement Learning from Mixed Feedback) in their training process to coordinate both ChatGPT-extracted data and real-world data from doctors, resulting in HuatuoGPT1 [24]. Chen et al., aiming to reduce the complexity of multi-stage training strategies, proposed a one-stage domain adaptation protocol where heterogeneous data from both traditional pre-training and supervised stages are unified into a simple instruction-output pair format, which they used to train HuatuoGPT2 [16]. Tian et al. developed ChiMedGPT, following a process that included pre-training, supervised fine-tuning (SFT), and reinforcement learning from human feedback (RLHF) [25]. Other researchers have attempted to improve the quality of training data. For example, Bao et al. constructed a high-quality SFT dataset using three strategies: utilizing medical knowledge graphs, reconstructing real-world dialogues, and incorporating human-guided preference rephrasing, significantly enhancing the model’s response capabilities in both single-turn and multi-turn consultation scenarios [17]. Luo et al. built a rich, multi-task bilingual dataset in both Chinese and English and trained the Taiyi model [18]. Although these medical LLMs have achieved considerable success in modern medicine, their performance in the TCM domain is significantly lacking.
@@ 2.2 LLMs in TCM Domain
Previous Chinese medical LLMs have mostly focused on modern medical systems. Unlike modern medicine, TCM emphasizes a holistic view and individualized treatment, focusing on balancing Yin-Yang, harmonizing Qi (气) and blood (血), and coordinating the internal organs. TCM diagnoses are conducted through the four diagnostic methods (inspection, listening/smelling, inquiry, and palpation) to understand the patient’s overall condition [19]. The theoretical foundation of syndrome differentiation and disease diagnosis in TCM is unique and fundamentally different from modern medicine. Existing medical LLMs are unable to grasp the principles of TCM, resulting in unsatisfactory performance.
Some studies have specifically aimed at improving the capabilities of LLMs in TCM. For example, Zhang et al. built a TCM-specific corpus and trained the Qibo model, which was applied to the pipeline of TCM consultations [20]. Yang et al. constructed a real-world, multi-turn Chinese medical dialogue dataset, CMtMedQA, and trained the Zhongjing model using continual pre-training, SFT, and RLHF, which enhanced the model’s ability in active inquiry and multi-turn understanding within the context of TCM [21]. Hua et al. focused on traditional Chinese medicine, constructing datasets for TCM pre-training, traditional Chinese patent medicine (TCPM) Q&A, and the spleen and stomach herbal prescription recommendation, and used these datasets to train the Lingdan model [22]. Although these efforts have made significant progress in TCM LLMs, these models still lack real-world diagnostic experience in TCM and perform poorly on real-world syndrome differentiation and disease diagnosis tasks.
To address the aforementioned challenges, we propose BianCang, a novel TCM LLM based on Qwen2/2.5. BianCang first undergoes continual pre-training to inject large amounts of TCM and medical knowledge, as well as real medical records, into the model. It is then fine-tuned using a carefully curated set of TCM-specific instructions to activate the model’s internal knowledge and align these instructions with the model’s internal knowledge. This training strategy ensures consistency in the model’s knowledge before and after SFT, significantly improving its performance in tasks such as syndrome differentiation and disease diagnosis in TCM.
@@ 3 Methods
The real hospital records used in this study were approved by the Ethics Committee of the Affiliated Hospital of Shandong University of Traditional Chinese Medicine (No. AF/SC-08/03.0). The construction of BianCang consists of two stages: continual pre-training and supervised fine-tuning. From the perspective of cognitive science, this two-stage framework—knowledge injection during pre-training and knowledge alignment during fine-tuning—closely mirrors the ACT-R [26] cognitive architecture. In ACT-R, declarative knowledge (i.e., factual information) is stored in long-term memory in the form of chunks, while procedural knowledge guides task-specific behavior. Pre-training resembles the accumulation of declarative chunks, whereas fine-tuning acts as a process of chunk activation and proceduralization, enabling the model to use its internal knowledge in a goal-directed manner. This theoretical framing helps justify why pre-training and fine-tuning must work together to achieve effective domain adaptation in large language models. The overall process is shown in Fig. 2.
Table 1: The statistics of the continual pre-training data.
@@
3.1 Continual Pre-training
@@
Recent research indicates that supervised fine-tuning (SFT) is not an effective mechanism of learning domain-specific world knowledge, but rather a process of aligning instructions with the knowledge already embedded within large language models [27]. Therefore, we should strive to inject extensive knowledge of traditional Chinese medicine into the foundational model during the pre-training phase [28].
Books serve as a high-quality source of data for training large language models [29, 30]. To construct a comprehensive corpus of Traditional Chinese Medicine (TCM) and general medical knowledge, we collected a substantial number of digitized books. These texts were processed using Optical Character Recognition (OCR), followed by a cleaning pipeline that applied heuristic rules to remove noisy characters and normalize text formatting. Manual sampling was performed to verify the quality of the processed content. For text segmentation, we employed the LlamaIndex [^3] https://github.com/run-llama/llama_index framework with a sliding-window strategy (i.e., overlapping chunking), using a chunk size of 1024 tokens to maintain contextual coherence.
In addition to textbook data, we incorporated open-source medical encyclopedias, scientific literature, and the ChatMed-TCM dataset [16, 31]. To further enrich the model’s TCM-specific knowledge, we integrated authoritative resources such as the Pharmacopoeia of the People’s Republic of China (Part I: Traditional Chinese Medicine) and commonly used TCM syndrome knowledge bases. These resources were manually inspected and cleaned on a file-by-file basis to ensure accuracy and consistency.
Recognizing that practical clinical experience is indispensable to TCM reasoning, we further included 12,259 real-world patient case records and 56,928 TCM syndrome differentiation and diagnosis records into the pre-training corpus. Lastly, to enhance the model’s performance on domain-specific assessments, we incorporated the training split of the Chinese Medical Benchmark (CMB) [32], which evaluates model capabilities across various real-world clinical scenarios.
@@
Besides TCM and medical corpora, we also incorporated general domain corpora into the pre-training data to prevent the model from overfocusing on specific domains, which could lead to catastrophic forgetting in general domains. Among them, COIG-CQIA is a high-quality, multi-domain instruction fine-tuning dataset aligned with human interaction behaviors [33]. APE-210K is a diverse collection of mathematical word problems [34]. Evol-Instruction-66K is a filtered and curated dataset for code generation [35]. We concatenated the instructions and responses from these three datasets as the pre-training data. Notably, all three general-domain datasets exhibit a certain level of logical reasoning, which we speculate could benefit tasks such as TCM syndrome differentiation, as it also requires logical reasoning skills.
@@
The data statistics for pre-training are shown in Table 1. Using this data, we performed continual pre-training on Qwen2/2.5-7B and Qwen2.5-14B [36, 37], eventually obtaining the BianCang TCM base model.
The negative log-likelihood (NLL) loss for continual pre-training of a language model can be defined as follows:
$\mathcal{L}_{\text{CPT}}(\theta)=$ (1)
$~~~~-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\log P_{\theta}(x_{t}^{(i)}\mid x_{1}^{(i)},\ldots,x_{t-1}^{(i)}),$
where $N$ denotes the number of training sequences (documents, sentences, etc.), $T_{i}$ denotes the length of the $i$-th sequence, $x_{t}^{(i)}$ denotes the token at position $t$ in the $i$-th sequence, $P_{\theta}(x_{t}^{(i)}\mid x_{1}^{(i)},\ldots,x_{t-1}^{(i)})$ denotes the probability of the token at position $t$ given the previous tokens, computed by the model with parameters $\theta$.
3.2 Supervised Fine-Tuning
Table 2: The statistics of the supervised fine-tuning.